ansform the equation 17𝑥

2 + 18𝑥𝑦 − 7𝑦

2 − 16𝑥 − 32𝑦 − 18 = 0 to one in 

which there is no term involving 𝑥, 𝑦.


1
Expert's answer
2021-01-20T17:47:19-0500

"17\ud835\udc65^2 + 18\ud835\udc65\ud835\udc66 \u2212 7\ud835\udc66^2 \u2212 16\ud835\udc65 \u2212 32\ud835\udc66 \u2212 18 = 0"


The rotation angle:

"tan2\theta=\frac{B}{A-C}"

where A, B and C are the corresponding coefficients from the equation:

"Ax^2+Bxy+Cy^2+Dx+Ey+F=0"

"tan2\theta=\frac{18}{1+7}=\frac{3}{4}=\frac{2tan\theta}{1-tan^2\theta}"

"tan\theta=\frac{1}{3}"


The required transformation equations needed to rotate the coordinate axes to eliminate the xy

term in the given equation for the conic section are:

"x=x'cos\theta-y'sin\theta"

"y=x'cos\theta+y'sin\theta"

"cos\theta=3/\sqrt{10}, sin\theta=1/\sqrt{10}"

"x=\frac{3x'-y'}{\sqrt{10}}, y=\frac{x'+3y'}{\sqrt{10}}"


Substituting these expressions for x and y into the original equation for the conic section and simplifying, we get:

"20x^{'2}-10y^{'2}-\frac{80}{\sqrt{10}}x'-\frac{80}{\sqrt{10}}y'-18=0"


Simplifying and completing the squares to obtain an equation for a hyperbola in standard form, we get:


"\frac{(x'-2/\sqrt{10})^2}{1/2}-\frac{(y'+4/\sqrt{10})^2}{1}=1"